OpenStudy (anonymous):

a circle d=12 and a center translated 18 units left and 7 units down from the orgin can some one check my answer. (x-y)^2 = (x+18)^2 x-y = x+18 -y = 18 y =-18 (y-k)^2 = (y+7)^2 y-k = y-7 -k = -7 y = 7

OpenStudy (anonymous):

(x + 18)^2 + (y + 7)^2 = 36 general formula is x^2 + y^2 = r^2

OpenStudy (anonymous):

thank